Chapter 4 (In-Class Assignment) Multiple Choice (CPA Adapted) 1. Parker, whose spouse died during the preceding year has not remarried. Parker maintains a home for a dependent child. What is Parkers most advantageous filing status? A. Single B. Head of Household C. Married Filing Separately D. Qualifying Widow(er) 2. Janet and Ted have two children, ages 8 & 10. Janets mother, Martha, resides in a nursing home nearby. Marthas only income is S1,500 a month in social security. Janet and Ted provide all of Marthas support. Does Martha qualify as Janet & Teds dependent? A. Yes, dependent child B. Yes, dependent relative C. No 3. Marie, an unmarried taxpayer, is 26 years old. This year, Marie earned S50,000 gross income. Her itemized deductions totaled S5,000. Marie maintained a home for her 12-year- old sister who qualifies as Maries dependent. Compute Maries 2018 taxable income. A. $45,000 B. $38,000 C. $32,000 D. $27,000 Tax Return Assignment Marc and Michelle Jones (SSNs: 111-22-3333 and 222-33-4444) are married and earned salaries this year of $64,000 and $20,000, respectively. In addition to their salaries, they received interest of S350 from municipal bonds and S500 from corporate bonds. Marc and Michelle have a 10- year-old son, Matthew (SSN: 999-88-7777), who lived with them throughout the entire year. Thus, Marc and Michelle are allowed to claim a $2,000 child tax credit for Matthew. Marc and Michelle paid S18,000 of expenses that qualify as itemized deductions and they had a total of S6,500 in federal income taxes withheld from their paychecks during the course of the year Required: Prepare Form 1040 for Marc and Michelle
Answer #1

1. Answer is option D

Qualifying widow(er)

As the spouse died, she is eligible to file as qualifying widow for two years. It is beneficial to file as qualifying widow as she gets all deductions of married filing jointly.

2. answer is option B

Yes, dependent relative

Martha qualifies as dependent relative of Janet and Ted because they provide all support to her. Moreover, she is Janet’s mother and thus it not compulsory for her stay with her to get qualify as dependent relative.

3. answer is option B

$38,000

Gross income – standard deduction = 50000-12000 = 38000

4. gross income = 84500 (64000+20000+500)

AGI

$84500 (64000+20000+500)

Standard deduction

24000

Itemized deductions

18000

Greater of standard deductions or itemized deductions

24000

Taxable income

60500

Income tax liability

6879 (1905+(12%*(60500-19050)))

Credits

(2000)

Prepayments

(6500)

Taxes refund

(1621)